OpenStudy (anonymous):

can anyone sketch f(x)= sec theta/2

Parth (parthkohli):

http://wolframalpha.com can do that.

OpenStudy (anonymous):

it would be great if you could show me how to arrive at the values

Parth (parthkohli):

You could just input this: f(x) = (secx)/2

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

go and download geobegra http://www.geogebra.org/cms/ it graphs everything

Parth (parthkohli):

Actually, the function should be:\[f(\theta) = {\sec \theta \over 2} \]OR\[ f(x) = {\sec x \over 2}\]

Parth (parthkohli):

\[f(x) = {\sec \theta \over 2} \]This does not make sense.

OpenStudy (anonymous):

place the sec before the fraction so you get sec*0/2

OpenStudy (anonymous):

make that sce*theta/2

OpenStudy (anonymous):

i'll try typing correctly again sec*theta/2

OpenStudy (unklerhaukus):

\[\sec\frac{\theta}{2}=\frac{1}{\cos\frac{\theta}{2}}\]

OpenStudy (unklerhaukus):